# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ 5.2. Motion of vertex positions
We derive bilinear equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
For $\Gamma\in\mathbb{R},y>0,r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(5.6) $\omega=\frac{\vartheta_{1}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)}{\vartheta_{3}\left(2r;iy\right)\vartheta_{1}\left(iv;iy\right)},$
$\delta_{1}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)^{3}},$
$\delta_{2}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{1}\left(iv;iy\right)^{2}\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)},$
$\delta_{3}=\frac{-iu_{2}}{s_{2}\vartheta_{3}(0;iy)\vartheta_{1}(iv;iy)}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right.$
$\left.+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right],$
$\epsilon_{1}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0;iy)}{\vartheta_{3}(0;iy)},$
$\epsilon_{2}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0;iy)^{2}}{\vartheta_{1}(iv;iy)^{2}},$
$\beta_{1}=\frac{1}{2}\left(\delta_{3}+\delta_{3}^{*}\right)+\frac{1}{2}i\Gamma\left(\delta_{2}-\delta_{2}^{*}\right),$
$\beta_{2}=\frac{1}{2}i\left(\delta_{3}-\delta_{3}^{*}\right)-\frac{1}{2}\Gamma\left(\delta_{2}+\delta_{2}^{*}\right).$
Then, it follows that
(5.7) $\omega^{*}=-\omega,\quad\delta_{1}^{*}=\delta_{1},\quad\epsilon_{1}^{*}=\epsilon_{1},\quad\epsilon_{2}^{*}=\epsilon_{2},\quad\beta_{1}^{*}=\beta_{1},\quad\beta_{2}^{*}=\beta_{2}.$
A proof of the following proposition is given in Appendix C.
@@
The $\tau$ functions (5.1) satisfy the following bilinear equations.
(5.8) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}+\omega G_{n}G_{n}^{*},$
(5.9) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\omega G_{n}G_{n}^{*},$
(5.10) $\delta_{2}f_{n}^{*}g_{n-1}-\delta_{2}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
(5.11) $\delta_{1}F_{n}F_{n}+\delta_{2}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
(5.12) $\left(\beta-\delta_{1}\right)F_{n}F_{n}+\left(\omega-\delta_{2}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
(5.13) $D_{t}H_{n}\cdot F_{n}=\left(\delta_{3}+i\Gamma\delta_{2}\right)f_{n}^{*}g_{n-1}+\left(\delta_{3}^{*}-i\Gamma\delta_{2}^{*}\right)f_{n-1}^{*}g_{n},$
(5.14) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\epsilon_{1}+C\right)F_{n}F_{n}+\epsilon_{2}G_{n}G_{n}^{*}.$
@@
By using these equations, we obtain the evolution equation of the vertex positions of the curve.
@@
(5.15) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta\left[\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right).\end{split}$
A proof of Theorem 5.6 is given in Appendix D. This is a deformation equation similar to the one presented in [16]. Starting from (5.15), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
@@
    (5.16) $\dot{\kappa}_{n}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
@@
The proofs of (i) and (ii) are given by the same calculation as in Section 5 of [16]. We prove (iii). From the deformation equation (5.15) and the discrete Frenet-Serret formula (2.6), we see that
(5.17) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta\beta_{1}\sin\lambda-2\beta\beta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
@@
Proposition 5.7 (iii) does not hold in general without the assumption on (5.27). We can construct a motion of a Kaleidocycle satisfying (5.15) such that the left-hand side of (5.27) depends both on $n$ and $t$.
@@
Set $C,\Gamma$ to be
(5.18) $\begin{split}\Gamma_{+}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)+\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{+}=&-\epsilon_{1}-i\Gamma_{+}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega},\\ \rm or\\ \Gamma_{-}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)-\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{-}=&-\epsilon_{1}-i\Gamma_{-}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega}.\end{split}$
Then, the curve constructed in Theorem 5.1 satisfies the following deformation equations.
(5.19) $\begin{split}&\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2},\end{split}$
where $\rho^{2}>0$ is given by
(5.20) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2}$
@@
(5.21) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2}$

### v3

@@ 5.2. Motion of vertex positions
We first derive equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
Lemma 5.3 ([12, §3]).
Using (3.2) and (3.3), we see that the following equations hold.
(5.41) $f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}=f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}=\beta_{1}F_{n}F_{n},$
(5.42) $g_{n}f_{n-1}-f_{n}g_{n-1}=G_{n}F_{n}.$
@@
The $\tau$ functions (5.1) satisfy the following equations.
(5.43) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta_{1}F_{n}F_{n}+\beta_{2}G_{n}G_{n}^{*},$
(5.44) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\beta_{2}G_{n}G_{n}^{*},$
(5.45) $\beta_{4}f_{n}^{*}g_{n-1}-\beta_{4}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
(5.46) $\beta_{3}F_{n}F_{n}+\beta_{4}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
(5.47) $\left(\beta_{1}-\beta_{3}\right)F_{n}F_{n}+\left(\beta_{2}-\beta_{4}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
(5.48) $D_{t}H_{n}\cdot F_{n}=\left(\beta_{5}+i\Gamma\beta_{4}\right)f_{n}^{*}g_{n-1}+\left(\beta_{5}^{*}-i\Gamma\beta_{4}^{*}\right)f_{n-1}^{*}g_{n},$
(5.49) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\beta_{6}+C\right)F_{n}F_{n}+\beta_{7}G_{n}G_{n}^{*}.$
where
(5.50) $\beta_{2}=\frac{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)}{\vartheta_{3}\left(2r\right)\vartheta_{1}\left(iv\right)},$
$\beta_{3}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)^{2}}{\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)^{3}},\quad\beta_{4}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)^{2}}{\vartheta_{1}\left(iv\right)^{2}\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)},$
$\beta_{5}=\frac{-iu_{1}}{\vartheta_{3}(0)\vartheta_{3}(2r)\vartheta_{1}(iv)^{2}}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right],$
$\beta_{6}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0)}{\vartheta_{3}(0)},\quad\beta_{7}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0)^{2}}{\vartheta_{1}(iv)^{2}}.$
Proof.
Putting
[…]
in (A.3) yields
(5.61) $\begin{split}&\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it+h\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\\ &=\vartheta_{4}\left(\mu_{n}-r+it\right)\vartheta_{3}\left(-\frac{1}{2}iv-r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right)\vartheta_{2}\left(\mu_{n-1}-r+it\right)\\ &-\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(-\frac{1}{2}iv-r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\vartheta_{4}\left(\mu_{n-1}-r+it\right),\end{split}$
After differentiating (5.61) by $h$ and putting $h=0$, we obtain (5.48) by rearrangement. Also, rearranging the equation with $h=0$ in (5.61) yields (5.45). Putting
[…]
in (A.2) yields
(5.72) $\begin{split}&\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it+h\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\\ &=-\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right)^{2}\frac{\vartheta_{1}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}+\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)^{2}\frac{\vartheta_{3}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}.\end{split}$
After differentiating (5.72) twice with $h$ and putting $h=0$, we obtain (5.49). ∎
Coefficients $\beta_{2}$ and $\beta_{4}$ in Proposition 5.4 satisfy the following relation.
Lemma 5.5.
The following relation holds.
(5.73) $\beta_{2}=\beta_{4}-\beta_{4}^{*}.$
Proof.
Putting
[…]
in (A.9) yields (5.73). ∎
Furthermore, to simplify the calculation of $\dot{\gamma_{n}}$, we define $P$ and $Q$ as follows
(5.79) $P_{n}=g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*},$
(5.80) $Q_{n}=g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}.$
Then the following expressions hold
(5.81) $D_{t}H_{n}\cdot F_{n}=\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*},$
(5.82) $\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}=T_{n}-\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054ptQ_{n}^{*}+Q_{n}\\ \vskip 5.69054pt\frac{1}{i}\left(Q_{n}^{*}-Q_{n}\right)\\ -2\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}+2\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\end{array}\right),$
(5.83) $\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\frac{1}{i}\left(P_{n}-P_{n}^{*}\right)\\ \vskip 5.69054ptP_{n}+P_{n}^{*}\\ -\frac{2}{i}\beta_{2}|G_{n}|^{2}\end{array}\right),$
where
(5.84) $\eta_{1}=\frac{1}{2}\left(\beta_{5}+\beta_{5}^{*}\right)+\frac{1}{2}i\Gamma\left(\beta_{4}-\beta_{4}^{*}\right),\quad\eta_{2}=\frac{1}{2}i\left(\beta_{5}-\beta_{5}^{*}\right)-\frac{1}{2}\Gamma\left(\beta_{4}+\beta_{4}^{*}\right).$
We have used (3.2) and (3.3) for (5.83). Then we have the following lemmas.
Lemma 5.6.
The following relations hold.
(5.85) $2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}=2\beta_{4}\beta_{5}^{*}+2\beta_{4}^{*}\beta_{5}+\beta_{7}=0,$
(5.86) $4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}+8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)-4\eta_{2}^{2}\beta_{2}^{2}-\beta_{7}^{2}=0.$
Proof.
By the definition of $\eta_{1}$ and $\eta_{2}$ and Lemma 5.5, we can see that the leftmost and middle part of (5.85) match. Putting
[…]
in (A.9) yields
(5.92) $\begin{split}&\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r+h\right)\\ &-\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r+h\right)\\ &=\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}\left(2h\right).\end{split}$
After differentiating (5.92) by $h$ and setting $h=0$, we obtain (5.85). From
(5.93) $\begin{split}&4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}+8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)-4\eta_{2}^{2}\beta_{2}^{2}-\beta_{7}^{2}\\ =&\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}\right\}\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}-\beta_{7}\right\},\end{split}$
we obtain (5.86). ∎
Lemma 5.7.
The following relations hold.
(5.94) $P_{n}P_{n}^{*}=F_{n}^{2}|G_{n}|^{2}+\beta_{2}^{2}|G_{n}|^{4},$
(5.95) $\begin{split}P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
(5.96) $\begin{split}&Q_{n}Q_{n}^{*}=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
Proof.
By definition of $P$, we see that
(5.97) $\begin{split}P_{n}P_{n}^{*}&=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\\ &=\left(f_{n}^{*}g_{n-1}^{*}-g_{n}^{*}f_{n-1}^{*}\right)\left(f_{n}g_{n-1}-g_{n}f_{n-1}\right)\\ &+\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right),\end{split}$
holds. Substituting (5.42) and (5.44) into (5.97) yields (5.94). By definition of $P$ and $Q$, we see that
[…]
holds. On the other hand, from (5.41) and (5.44), the following holds.
[…]
Using (5.99), we obtain
(5.100) $\begin{split}&P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}\\ &=f_{n}f_{n-1}^{*}f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}f_{n}^{*}f_{n-1}+g_{n}g_{n-1}^{*}g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}g_{n}^{*}g_{n-1}\\ &=\left(f_{n}f_{n-1}^{*}\right)^{2}-\left(f_{n}^{*}f_{n-1}\right)^{2}+\left(g_{n}g_{n-1}^{*}\right)^{2}-\left(g_{n}^{*}g_{n-1}\right)^{2}\\ &=\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\\ &+\left(g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right).\end{split}$
Substituting (5.44), (5.46) and (5.47) into (5.100) gives
(5.101) $\begin{split}&P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}\\ &=\beta_{2}|G_{n}|^{2}\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}-g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\\ &=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
thus (5.95). Also from (5.42), (5.46) and (5.47), we see that
(5.102) $\begin{split}&Q_{n}Q_{n}^{*}-F_{n}^{2}|G_{n}|^{2}\\ &=\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}-f_{n}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}^{*}-f_{n}^{*}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right)\\ &=\left(2\beta_{3}F_{n}^{2}+\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\right)\left(2\left(\beta_{1}-\beta_{3}\right)F_{n}^{2}-\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\right)\\ &=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)F_{n}^{2}|G_{n}|^{2}-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4},\end{split}$
holds. Thus we get (5.96). ∎
@@
By using Proposition 5.4, Lemma 5.5 Lemma 5.6 and Lemma 5.7, we obtain the evolution equation of the vertex positions of the curve.
@@
(5.103) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta_{1}\left[\eta_{1}T_{n}-\eta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)-\beta_{6}-C\end{array}\right).\end{split}$
Proof.
By differentiating $\gamma$ with respect to $t$, we get
(5.104) $\begin{split}\dot{\gamma}_{n}(t)&=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle D_{t}H_{n}\cdot F_{n}+D_{t}H_{n}^{*}\cdot F_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(D_{t}H_{n}\cdot F_{n}-D_{t}H_{n}^{*}\cdot F_{n}\right)\\ -D_{t}D_{z}F_{n}\cdot F_{n}\end{array}\right)\\ &=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -\left(\beta_{6}+C\right)F_{n}^{2}-\beta_{7}|G_{n}|^{2}\end{array}\right).\end{split}$
On the other hand, from (5.82) and (5.83), we see
(5.105) $\begin{split}&\eta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}+2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}+2i\eta_{2}\beta_{2}|G_{n}|^{2}\end{array}\right)\end{split}$
holds. From (5.85), (5.105) can be rewritten as follows
(5.106) $\begin{split}&\eta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}-\beta_{7}|G_{n}|^{2}\end{array}\right).\end{split}$
Thus (5.103) is obtained. ∎
This is a deformation equation similar to the one presented in [18]. Starting from (5.103), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
@@
    (5.107) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
@@
The proofs of (i) and (ii) follow from the same calculation as in Section 5 of [18]. We prove (iii). From the deformation equation (5.103) and the discrete Frenet-Serret formula (2.5), we see that
(5.108) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta_{1}\eta_{1}\sin\lambda-2\beta_{1}\eta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
@@
Proposition 5.9 (iii) does not hold in general without the assumption on (5.136). Indeed, there is a motion of a Kaleidocycle satisfying (5.103) such that the left-hand side of (5.136) depends both on $n$ and $t$.
@@
Set $C$ and $\Gamma$ to be
(5.109) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0,\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.\end{split}$
Then, the curve constructed in Theorem 5.1 deforms according to the equation (3.24). The magnitude of $\dot{\gamma}_{n}$ is a constant and is given by
(5.110) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2},$
@@
(5.111) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2},$